OpenStudy (anonymous):

For this question, how is the answer 3?: Determine the slope of the tangent to the curve y = (3x)(sinx) at the point with x-coordinate π/2.

OpenStudy (anonymous):

i think i got it

OpenStudy (anonymous):

i found the derivative to be 3cosx^2 + 3sin x and subbed in π/2. then i got 3.079 as my answer

OpenStudy (anonymous):

pi/2

OpenStudy (displayerror):

How'd you get 3cosx^2? The answer should come out to be exactly 3. \[f(x) = 3x\] \[f'(x) = 3\] \[g(x) = \sin{x}\] \[g'(x) = \cos{x}\]

OpenStudy (anonymous):

i got that too. then i did this: y'=(3x)(cosx)+(3)(sinx) y'=3cosx^2 + 3sinx

OpenStudy (displayerror):

You can't multiply the \(x\) in \(3x\) with the \(x\) in \(\cos{x}\). It's two separate functions, \(3x\) and \(\cos{x}\).

OpenStudy (anonymous):

ohh but when I don't, I get 4.79 as my answer

OpenStudy (displayerror):

The derivative should be \[f'(x) = 3 \sin{x} + 3x\cos{x}\] Evaluate with x = pi/2: \[f'(\frac{\pi}{2}) = 3 \sin{\frac{\pi}{2}} + \frac{3\pi}{2}\cos{\frac{\pi}{2}}\] What does \(\sin{(\pi/2})\) equal? What does \(\cos{(\pi/2})\) equal?

OpenStudy (anonymous):

sin(pi/2) = 0.02741 cos(pi/2) = 0.9996

OpenStudy (displayerror):

Your calculator is in degree mode. Change it to radian mode and re-evaluate.

OpenStudy (anonymous):

Oh! now I'm getting 3!

OpenStudy (anonymous):

thanks a lot!